# CinePile: A Long Video Question Answering Dataset and Benchmark

### v1

@@ Abstract
Current datasets for long-form video understanding often fall short of providing genuine long-form comprehension challenges, as many tasks derived from these datasets can be successfully tackled by analyzing just one or a few random frames from a video. To address this issue, we present a novel dataset and benchmark, CinePile, specifically designed for authentic long-form video understanding. This paper details our innovative approach for creating a question-answer dataset, utilizing advanced LLMs with human-in-the-loop and building upon human-generated raw data. Our comprehensive dataset comprises 305,000 multiple-choice questions (MCQs), covering various visual and multimodal aspects, including temporal comprehension, understanding human-object interactions, and reasoning about events or actions within a scene. Additionally, we evaluate recent video-centric LLMs, both open-source and proprietary, on the test split of our dataset. The findings reveal that even state-of-the-art video-centric LLMs significantly lag behind human performance in these tasks, highlighting the complexity and challenge inherent in video understanding.
footnotetext:
Equal contribution. Correspondence: {ruchitr,gowthami}@umd.edu
@@ 1 Introduction
In this paper, we introduce CinePile, a large-scale dataset consisting of $\sim$ 305k question-answer pairs from 9396 videos, split into train and test sets. Our dataset emphasizes question diversity, and topics span temporal understanding, perceptual analysis, complex reasoning, and more. It also emphasizes question difficulty, with humans exceeding the best commercial vision/omni models by approximately 26%, and exceeding open source video understanding models by 70%.
@@
CinePile addresses several weaknesses of existing video understanding datasets. First, the large size of CinePile enables it to serve as both an instruction-tuning dataset and an evaluation benchmark. We believe the ability to do instruction tuning for video at a large scale can bridge the gap between the open-source and commercial video understanding models. Also, the question diversity in CinePile makes it a more comprehensive measure of model performance than existing benchmarks. Unlike existing metrics (see Section 6 for detailed comparisons), CinePile does not over-emphasize on purely visual questions (e.g., What color is the car?), or on classification questions (e.g., What genre is the video?) that do not require temporal understanding. Rather, CinePile is comprehensive with diverse questions about vision, temporal, and narrative reasoning with a breakdown of question types to help developers identify blind spots in their models.
The large size of CinePile is made possible by our novel pipeline for automated question generation and verification using large language models. Our method leverages large existing sets of audio descriptions that have been created to assist the vision impaired. We transcribe these audio descriptions and align them with publicly available movie video clips from YouTube. Using this detailed human description of scenes, powerful LLMs are able to create complex and difficult questions about the whole video without using explicit video input. At test time, video-centric models must answer these questions from only the dialogue and raw video, and will not have access to the hand-written descriptions used to build the questions. Dataset and artifacts are available at this https://ruchitrawal.github.io/cinepile.
@@ 2 Creating a long video understanding benchmark
Our dataset curation process has four primary components 1) Collection of raw video and related data. 2) Generation of question templates. 3) Automated construction of the question-answer dataset using video and templates, and 4) Application of a filtering pipeline to remove malformed questions.
@@ 2.1 Data collection and consolidation
We obtain clips from English-language films from the YouTube channel MovieClips [^1] https://www.youtube.com/@MOVIECLIPS. This channel hosts self-contained clips, each encapsulating a major plot point, facilitating the creation of a dataset focused on understanding and reasoning. Next, we collected Audio Descriptions from AudioVault [^2] https://audiovault.net/movies. Lastly, we collect movie information for each scene, such as genre, actors, and main plot from IMDB [^3] https://www.imdb.com/.
@@
Scene localization in AD. The video clips we have gathered are typically 2-3 minutes long, while Audio Descriptions (ADs) cover entire movies. To align descriptions with video, we transcribe the audio from both the movie clip and the whole movie AD file using an Automatic Speech Recognition (ASR) system WhisperX [4], an enhanced version of Whisper [22] designed to offer quicker inference and more precise word-level timestamps. We then embed the first 3 and last 3 lines of the text transcription of a YouTube movie clip using a sentence embedding model, WhereIsAI/UAE-Large-V1. We similarly embed all the sentences in the corresponding movie AD file. We then localize the YouTube clip within the AD file via the rolling window algorithm. We then extract all AD data that lies between the matched start and end of the movie clip embeddings. This localized text contains both the visual elements and the dialogue for the given YouTube clip. This serves as a base text for creating the QA dataset which we will discuss in the rest of the paper.
For the rest of the paper, we will refer to the human-written description of the scene as “visual description” and the speaking or dialogue part of the video as “dialogue”. When combined, we will refer to both data sources as “scene-text-annotation”.
Sentence classification. When we transcribe an AD file, the text contains a human’s visual descriptions and the movie’s dialogue. However, the transcription model does not label whether a given sentence belongs to a visual description or a dialogue. Since we planned to create a few questions solely on the visual components of the video, the distinction is important to us. To categorize each sentence as either visual or dialogue, we fine-tuned a BERT-Base model [7] using annotations from the MAD dataset [25], which contains labels indicating whether a sentence is a dialogue or a visual description. We applied a binary classification head for this task. For training the classification model, we split the MAD dataset annotations into an 80-20 training-evaluation split. The model achieves 96% accuracy on eval split after 3 epoch training. Qualitatively, we observed that the model accurately classifies sentences in the data we curated, distinguishing effectively between dialogue and visual description content.
Figure 2: Question template generation pipeline: We begin by substituting the first names in human-written source questions and then cluster them. We then feed a selection of questions from each cluster into GPT-4 which in turn outputs “question templates” which are used in the next stage of dataset creation. See Section 2.2 for more details.
@@
Figure 3: Extracting templates from human-generated questions. We share 10 samples from each question cluster, and prompt an LLM to create a few templates and a prototypical question. See Fig. 2 and Section 2.2 for more details.
@@ 2.2 Automated Question Templates
Many prominent video question-answering benchmarks were written by human annotators. The question-answer pairs are typically curated in one of two ways: 1) Human annotators are given complete freedom to ask questions about a given scene [28] 2) They are asked to focus on specific aspects and are trained or provided with examples of questions, encouraging them to write more questions in a similar style [37, 13, 11, 21]. For instance, in the Perception Test Benchmark [21], annotators are directed to concentrate on temporal or spatial aspects, while for the Next-QA dataset [37], annotators mainly focused on temporal and causal action reasoning questions.
@@
During early experimentation, we found that providing a range of templates along with scene-text-annotation to an LLM helped it create more detailed, diverse, and well-formed questions, so we decided to use a template-based approach for question generation. Rather than confining questions to a few hand-curated themes, we propose a pipeline to create question templates naturally on top of human-generated questions. The automated question template generation pipeline is illustrated in Fig. 2.
@@
Our starting point is approximately 30,000 human-curated questions from the MovieQA [28], TVQA [11], and Perception Test [21] datasets. We cluster these questions, select a few representatives per cluster, and then use GPT-4 to discern the underlying themes and write a prompt. First, we preprocess the questions by replacing first names and entities with pronouns, as BERT [24] embeddings over-index on proper nouns, hence the resultant clusters end up with shared names rather than themes. For instance, ‘Why is Rachel hiding in the bedroom?’ is altered to ‘Why is she hiding in the bedroom?’. We used GPT-3.5 to do this replacement, as it handled noun replacement better than many open-source and commercial alternatives. The modified questions are then embedded using WhereIsAI/UAE-Large-V1, a semantic textual similarity model which is a top performer on the MTEB leaderboard [^4] https://huggingface.co/spaces/mteb/leaderboard. When the first names were replaced, we observed significant repetition among questions, which prompted us to duplicate them, ultimately resulting in 17,575 unique questions. We then perform k-means clustering to categorize the questions into distinct clusters. We experimented with different values of $k=10,50,100$. Qualitatively, we found $k=50$ to be an optimal number of clusters where the clusters are diverse and at the same time clusters are not too specific. For example, we see a ‘high-school dance’ specific question cluster when $k=100$, and these questions are merged into an ‘event’ cluster when we reduce the number of clusters to 50. The Perception Test questions are less diverse as human annotators were restricted to creating questions based on a small number of themes, so we used $k=20$ for this set.
The number of questions in each cluster varied, with counts ranging from 60 to 450. We selected 10 random questions from each, and used them to prompt GPT-4 to create relevant question templates, as illustrated in Fig. 3. We also ablated the scenario of taking the closest 10 questions to the cluster center, we however qualitatively observed that random questions produced more general/ higher quality templates.

### v3

@@ Abstract
Current datasets for long-form video understanding often fall short of providing genuine long-form comprehension challenges, as many tasks derived from these datasets can be successfully tackled by analyzing just one or a few random frames from a video. To address this issue, we present a novel dataset and benchmark, CinePile, specifically designed for authentic long-form video understanding. This paper details our innovative approach for creating a question-answer dataset, utilizing advanced LLMs with human-in-the-loop and building upon human-generated raw data. Our comprehensive dataset comprises 305,000 multiple-choice questions (MCQs), covering various visual and multimodal aspects, including temporal comprehension, understanding human-object interactions, and reasoning about events or actions within a scene. Additionally, we fine-tuned open-source Video-LLMs on the training split and evaluated both open-source and proprietary video-centric LLMs on the test split of our dataset. The findings indicate that although current models underperform compared to humans, fine-tuning these models can lead to significant improvements in their performance.
@@ 1 Introduction
In this paper, we introduce CinePile, a large-scale dataset consisting of $\sim$ 305k question-answer pairs from 9396 videos, split into train and test sets. Our dataset emphasizes question diversity, and topics span temporal understanding, perceptual analysis, complex reasoning, and more. It also emphasizes question difficulty, with humans exceeding the best commercial vision/omni models by approximately 25%, and exceeding open source video understanding models by 37%.
@@
CinePile addresses several weaknesses of existing video understanding datasets. First, the large size of CinePile enables it to serve as both an instruction-tuning dataset and an evaluation benchmark. We believe the ability to do instruction tuning for video at a large scale can bridge the gap between the open-source and commercial video understanding models. Also, the question diversity in CinePile makes it a more comprehensive measure of model performance than existing benchmarks. Unlike existing datasets, CinePile does not over-emphasize on purely visual questions (e.g., What color is the car?), or on classification questions (e.g., What genre is the video?) that do not require temporal understanding. Rather, CinePile is comprehensive with diverse questions about vision, temporal, and narrative reasoning with a breakdown of question types to help developers identify blind spots in their models.
The large size of CinePile is made possible by our novel pipeline for automated question generation and verification using large language models. Our method leverages large existing sets of audio descriptions that have been created to assist the vision impaired. We transcribe these audio descriptions and align them with publicly available movie video clips from YouTube. Using this detailed human description of scenes, powerful LLMs are able to create complex and difficult questions about the whole video without using explicit video input. At test time, video-centric models must answer these questions from only the dialogue and raw video, and will not have access to the hand-written descriptions used to build the questions. We release the prompts for generating the question answers, the code for model evaluation, and the dataset splits in the Appendix.
@@ 2 Creating a long video understanding benchmark
Our dataset curation process has four primary components 1) Collection of raw video and related data. 2) Generation of question templates. 3) Automated construction of the Q&A dataset using video and templates, and 4) Application of a refinement pipeline to improve or discard malformed Q&A pairs.
@@ 2.1 Data collection and consolidation
We obtain clips from English-language films from the YouTube channel MovieClips [^1] https://www.youtube.com/@MOVIECLIPS. This channel hosts self-contained clips, each encapsulating a major plot point, facilitating the creation of a dataset focused on understanding and reasoning. Next, we collected Audio Descriptions from AudioVault [^2] https://audiovault.net/movies.
@@
Scene localization in AD. The video clips we have gathered are typically 2-3 minutes long, while Audio Descriptions (ADs) cover entire movies. To align descriptions with video, we transcribe the audio from both the movie clip and the whole movie AD file using an Automatic Speech Recognition (ASR) system WhisperX (Bain et al., 2023), an enhanced version of Whisper (Radford et al., 2023) designed to offer quicker inference and more precise word-level timestamps. We then embed the first 3 and last 3 lines of the text transcription of a YouTube movie clip using a sentence embedding model, WhereIsAI/UAE-Large-V1. We similarly embed all the sentences in the corresponding movie AD file. We then localize the YouTube clip within the AD file via the rolling window algorithm. We then extract all AD data that lies between the matched start and end of the movie clip embeddings. This localized text contains both the visual elements and the dialogue for the given YouTube clip. This serves as a base text for creating the QA dataset For the rest of the paper, we will refer to the human-written description of the scene as “visual description” and the speaking or dialogue part of the video as “dialogue”. When combined, we will refer to both data sources as “scene-text-annotation”.
Sentence classification. When we transcribe an AD file, the text contains a human’s visual descriptions and the movie’s dialogue. However, the transcription model does not label whether a given sentence belongs to a visual description or a dialogue. Since we planned to create a few questions solely on the visual components of the video, the distinction is important to us. To categorize each sentence as either visual or dialogue, we fine-tuned a BERT-Base model (Devlin et al., 2018) using annotations from the MAD dataset (Soldan et al., 2022), which contains labels indicating whether a sentence is a dialogue or a visual description. We applied a binary classification head for this task. For training the classification model, we split the MAD dataset annotations into an 80-20 training-evaluation split. The model achieves 96% accuracy on eval split after 3 epoch training. Qualitatively, we observed that the model accurately classifies sentences in the data we curated, distinguishing effectively between dialogue and visual description content.
22footnotetext: Icons in the figures are sourced from Flaticon.
Figure 2: Question template generation pipeline: We begin by substituting the first names in human-written source questions and then cluster them. We then feed a selection of questions from each cluster into GPT-4, which outputs “question templates” used in the next stage of dataset creation. See Section 2.2 for more details.
@@
Figure 3: Extracting templates from human-generated questions. We share 10 questions from each cluster, and prompt an LLM to create a few templates and a prototypical question. See Fig. 2 and Section 2.2 for details.
@@ 2.2 Automated Question Templates
Many prominent video question-answering benchmarks were written by human annotators. The question-answer pairs are typically curated in one of two ways: 1) Human annotators are given complete freedom to ask questions about a given scene (Tapaswi et al., 2016) 2) They are asked to focus on specific aspects and are trained or provided with examples of questions, encouraging them to write more questions in a similar style (Xiao et al., 2021; Li et al., 2020; Lei et al., 2018; Patraucean et al., 2024). For instance, in the Perception Test Benchmark (Patraucean et al., 2024), annotators are directed to concentrate on temporal or spatial aspects, while for the Next-QA dataset (Xiao et al., 2021), annotators mainly focused on temporal and causal action reasoning questions.
Figure 4: Automated QA Generation and Filtering. Begins with a set of automated templates and scenes. Filter out the templates relevant to each scene. Next, pass these templates along with the annotated-scene-text to GPT-4, which is then used to create multiple-choice questions (MCQs). The generated MCQs are then subjected to numerous filters to curate the final dataset. For more detailed information, refer to Section 2.3 and Section 2.4
@@
During early experiments, we found that giving a range of templates and scene-text-annotation to an LLM helped create more detailed, diverse, and well-formed questions. Thus, we adopted a template-based approach for question generation. Instead of limiting questions to a few hand-curated themes, we propose a pipeline to create templates from human-generated questions (shown in Fig. 2).
@@
Our starting point is approximately 30,000 human-curated questions from the MovieQA (Tapaswi et al., 2016), TVQA (Lei et al., 2018), and Perception Test (Patraucean et al., 2024) datasets. We cluster these questions, select a few representatives per cluster, and then use GPT-4 to discern the underlying themes and write a prompt. First, we preprocess the questions by replacing first names and entities with pronouns, as BERT (Reimers & Gurevych, 2019) embeddings over-index on proper nouns, hence the resultant clusters end up with shared names rather than themes. For instance, ‘Why is Rachel hiding in the bedroom?’ is altered to ‘Why is she hiding in the bedroom?’. We used GPT-3.5 to do this replacement, as it handled noun replacement better than many open-source and commercial alternatives. The modified questions are then embedded using WhereIsAI/UAE-Large-V1, a semantic textual similarity model which is a top performer on the MTEB leaderboard [^3] https://huggingface.co/spaces/mteb/leaderboard. When the first names were replaced, we observed significant repetition among questions, which prompted us to duplicate them, ultimately resulting in 17,575 unique questions. We then perform k-means clustering to categorize the questions into distinct clusters. We experimented with different values of $k=10,50,100$. Qualitatively, we found $k=50$ to be an optimal number of clusters where the clusters are diverse and at the same time clusters are not too specific. For example, we see a ‘high-school dance’ cluster when $k=100$, and these questions are merged into an ‘event’ cluster when we reduce $k$ to 50. The Perception Test questions are less diverse as human annotators were restricted to creating questions based on a small number of themes, so we used $k=20$ for this set. The number of questions in each cluster ranges from 60 to 450. We selected 10 random questions from each, and used them to prompt GPT-4 to create relevant question templates, as illustrated in Fig. 3. We did ablations by selecting the closest 10 questions to the cluster center, however qualitatively observed that random questions produced more general/higher quality templates.